# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v2

@@ 3. Unsupervised Video Anomaly Detection
UVAD follows the general setup of semantic anomaly detection tasks, where only easily collectible regular instances are used to train models describing the normality of videos, detecting anomalies by measuring deviations between test samples and the learned model (Cheng et al., 2024a, b; Wang et al., 2021b). From traditional machine learning to deep representation learning (Yang et al., 2024a; Wang et al., 2024b; Yang et al., 2023b), UVAD has undergone multiple advancements in feature extraction and normality learning, leading VAD to become a key issue in the AD and CV communities. With its close ties to the AD community and its long-standing development history, UVAD has long been regarded by researchers as the mainstream research route of NSVAD algorithms. As a result, existing surveys (Nayak et al., 2021; Ramachandra et al., 2020) typically focus on reviewing UVAD literature, lacking in-depth discussion of emerging weakly supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Zaheer et al., 2022; Pang et al., 2020) routes, and overlooking novel tasks such as open-set and open-word detection. This article not only systematically reviews the latest developments in these new routes and tasks but also provides foundational knowledge and timely and classic methods of UVAD in this section.
@@
Based on the means of normality learning and deviation calculation principles, existing deep learning methods are generally divided into three categories: distance-based (Saligrama and Chen, 2012; Dalal et al., 2006; Dalal and Triggs, 2005), probability-based (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), and reconstruction-based (Liu et al., 2024b, 2023f; Bao et al., 2022; Luo et al., 2017). From our perspective, distance-based UVAD methods are a more general form of probability-based and reconstruction-based methods because probability deviation and reconstruction error calculation are fundamentally just different distance measurement methods. Specifically, distance-based methods include using single-classifier learning of video sequence spatial-temporal representations in the deep feature space, such as OC-SVM and OC-NN (Ionescu et al., 2019; Sun et al., 2019). The drawback of such methods is that the trained models cannot be incrementally compatible with new data, leading to the need to retrain classifiers from scratch when new data is generated, such as in scenarios involving scene transitions. Another type of distance-based method is to use Gaussian mixture models (Wu et al., 2010) to model video feature normal vectors and measure deviations using Mahalanobis distance. In contrast, probability-based methods attempt to map the spatial-temporal representations of regular events into a probability framework and discriminate anomaly instances by measuring differences in probability distributions. Such methods tend to use traditional models such as Markov random fields (Benezeth et al., 2009) to build probability space. Deep learning-based attempts have encountered significant increases in computational costs and are noticeably slower in inference speeds.
@@
In fact, the most prevailing deep UVAD approach is reconstruction-based (Ramachandra et al., 2020). On the one hand, reconstruction-based methods (Liu et al., 2018b; Park et al., 2020; Liu et al., 2024b) aim to train models to represent the general spatial-temporal patterns of regular events through self-supervised proxy tasks, benefiting from advances in video self-supervised learning (Jing and Tian, 2020) and deep neural networks. On the other hand, such methods avoid complex mathematical computations, which are easy to implement and exhibit excellent performance, thus being widely praised by existing researchers. The premise of reconstruction-based methods is that generative models trained on massive normal samples can effectively infer the spatial-temporal patterns of regular events. For anomalous events, the performance of proxy tasks will significantly decrease, and the resulting error can be used as a quantitative basis for measuring deviations to calculate anomaly scores. Common generative models include deep autoencoders (Hasan et al., 2016; Chong and Tay, 2017; Liu et al., 2022a, c; Park et al., 2020), variational autoencoders (Fan et al., 2020; Yan et al., 2018), and generative adversarial networks (Chen et al., 2021; Huang et al., 2021; Nguyen et al., 2020; Li et al., 2022; Cai et al., 2021a; Zhang et al., 2022). Proxy tasks include reconstructing input sequences and predicting future frames, which belong to pixel-level image generation. However, most reconstruction-based methods only calculate frame-level errors as anomaly scores without performing spatial localization, as the spatial contribution of anomalies is typically not significant. Spatial localization can only serve as a quantitative visualization result.
@@ 3.1. Taxonomy and Advances
UVADGNLLPM Single-task: (Medel and Savakis, 2016), (Hasan et al., 2016), (Hu et al., 2016), (Smeureanu et al., 2017),(Luo et al., 2017),(Liu et al., 2018b),(Luo et al., 2021),(Singh and Pankajakshan, 2018),(Li et al., 2019), (Zhou et al., 2019),(Nawaratne et al., 2019),(Ye et al., 2019),(Gong et al., 2019),(Park et al., 2020),(Luo et al., 2019),(Deepak et al., 2021),(Dong et al., 2020),(Yu et al., 2021),(Zhao et al., 2022),(Chen et al., 2020),(Wang et al., 2021a),(Nayak et al., 2023) Multi-tasks: (Chang et al., 2020),(Chang et al., 2021),(Li et al., 2021a),(Cai et al., 2021b),(Liu et al., 2022a),(Liu et al., 2022c),(Zhao et al., 2017),(Xu et al., 2017),(Nguyen and Meunier, 2019),(Ravanbakhsh et al., 2019),(Zaheer et al., 2020) ST Patch: (Mahadevan et al., 2010),(Zhou et al., 2016),(Christiansen et al., 2016),(Sabokrou et al., 2017),(Wu et al., 2019),(Li et al., 2020),(Liu et al., 2021a),(Wang et al., 2018),(Roshtkhari and Levine, 2013) Object-level: (Georgescu et al., 2021b),(Georgescu et al., 2021a),(Hinami et al., 2017),(Ko and Sim, 2018),(Ionescu et al., 2019),(Yu et al., 2020),(Doshi and Yilmaz, 2021),(Liu et al., 2021b),(Chen et al., 2022),(Bao et al., 2022),(Liu et al., 2023a)
@@
GNL utilizes the entire video sequence as input, often requiring no additional preprocessing such as spatial-temporal cube partitioning or foreground object extraction, and employs end-to-end deep neural networks to directly learn video ecologies (Huang et al., 2022; Gong et al., 2019; Park et al., 2020; Chang et al., 2020; Liu et al., 2024b, 2023e). Over the years, researchers have believed that videos possess two informational dimensions, namely spatial and temporal, corresponding to appearance and motion, requiring different approaches to capture their normality. Therefore, within our UVAD taxonomy, GNL is further subdivided into Single-Proxy Task (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020) and Multi-Proxy Task methods (Chang et al., 2020; Liu et al., 2022c, 2024b). In contrast, LPM methods argue that video data contain a plethora of redundant information tantamount to clues related to normality. Thus, they opt to use spatial-temporal cubes (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a) or foreground objects (Bao et al., 2022; Liu et al., 2023a) containing dense and effective information as network inputs instead of the entire video sequence, focusing on prototype feature learning of local patch. We classify such methods into spatial-temporal Patch-based and Object-driven methods.
@@ 3.1.1. Global Normality Learning
The earliest methods did not distinguish between spatial and temporal information, typically using only RGB videos as input and employing a single self-supervised proxy task (e.g., sequence reconstruction or future frame prediction (Zhao et al., 2024)) to broadly learn spatial-temporal normality. Generally, methods based on a single proxy task focus on designing more efficient single-stream end-to-end deep structures. Recent efforts include the introduction of 3D convolutional networks (Gong et al., 2019) and convolutional long short-term memory networks (Luo et al., 2017) to enhance the representation capability of spatial-temporal features. Subsequent researchers discovered that spatial and temporal information have different characteristics, with the former focusing on local pixel inference and the latter on modeling global dynamics. Moreover, the addition of proxy tasks such as reconstruction and prediction losses often brings additional performance gains without significantly increasing training costs. Thus, they proposed introducing additional proxy tasks within the GNL framework. In addition to the separation of spatial-temporal normality learning (Liu et al., 2022c; Chang et al., 2021) and simple proxy task stacking (Zhao et al., 2017; Chen et al., 2020), recent work has also explored novel tasks such as appearance-motion consistency (Cai et al., 2021b), spatial-temporal coherence (Cheng et al., 2023a; Ning et al., 2024), and correlation (Zhao et al., 2022).
@@ 4. Weakly-supervised Video Anomaly Detection
Inspired by multiple instance learning (MIL) (Herrera et al., 2016), WsVAD organizes videos into bags containing several instances. All segment instances from the same bag share a video-level label (Huang et al., 2024a). Here, $Y=0$ indicates negative bags while $Y=1$ indicates positive ones containing at least one anomalous instance. WsVAD strikes a balance between performance and data preparation cost, bridging the gap between UVAD with unreliable results and supervised learning with fine labels. Specifically, unlike unsupervised methods that train models only on regular events and simply consider all unseen samples as anomalies, WsVAD is trained with positive samples, enabling it to effectively understand the inherent differences between normal and anomalous instances. Consequently, existing research indicates that WsVAD yields more reliable results and outperforms UVAD schemes with lower false alarm rates in real-world tasks, such as crime behavior identification and traffic accident detection (Liu et al., 2023d). Annotating long video sequences frame by frame for supervised learning is often impractical. For instance, the training set of the UCF-Crime dataset includes 1,610 long videos containing 13,741,393 frames in total. In contrast, WsVAD only requires coarse semantic video-level annotations, with each video needing only a discrete binary label. Various types of anomalies are roughly labeled as 1, requiring minimal expert involvement in data annotation and significantly reducing labor costs compared to fine-grained labeling. Thus, although WsVAD no longer adheres to the open-world task setting of the AD community and can only detect specific anomalies, it has garnered widespread attention in recent years due to its stable performance and reliable results. Subsequent works typically adopt the MIL regression task setting proposed by Sultani et al. (Sultani et al., 2018) and utilize their concurrently open-sourced UCF-Crime dataset as a benchmark. To further validate the generalization ability of weakly-supervised models, some researchers relocate positive samples from the test set of UVAD datasets to the training sets and provide video-level labels, proposing reconfigured datasets such as UCSD Ped2 and ShanghaiTech Weakly for WsVAD validation.
@@ 4.1. Taxonomy and Advances
The superior performance of WsVAD on real-world videos has inspired researchers to develop VAD models tailored to the complexities of real-life scenarios. In 2020, Wu et al. (Wu et al., 2020) extended VAD research from single-modal video pattern analysis to multimodal learning to leverage heterogeneous data in real-world scenes for enhanced anomaly event detection. Their collected XD-violence dataset is the first multimodal VAD benchmark, comprising RGB images and audio modalities, focusing on violence behavior detection in complex scenes. This work not only provides the first multimodal VAD dataset and solution but also motivates researchers in the community to explore anomaly clues from multimodal data such as audio and text, leading to a new wave of VAD development: multimodal WsVAD. In this section, we comprehensively review the latest advancements in both single-modal and multimodal WsVAD, and illustrate the foundational knowledge and specific implementations required for WsVAD using the Multiple Instance Regression (MIR) framework and the Local-Global Network (HL-Net) as examples.
@@ 4.1.1. Unimodal Methods
Unimodal WsVAD (UWsVAD) methods focus on extracting anomaly-related cues from RGB image sequences (Liu et al., 2022d). These methods generally follow a three-step process: 1) preprocessing videos into several non-overlapping segments, 2) extracting spatial-temporal features using models like Convolutional 3D Networks (Tran et al., 2015) and Inflated 3D Networks (Carreira and Zisserman, 2017), and 3) computing anomaly scores using a multi-instance ranking loss to differentiate between normal and anomalous instances. The Multi-Instance Ranking (MIR) framework (Sultani et al., 2018) first introduced MIL to WsVAD with a focus on predicting higher anomaly scores for anomalous segments while minimizing score fluctuations. For a more detailed description of the existing UWsVAD methods, please refer to Section 1.2.1 of Supplement. Moreover, given that VsVAD datasets are typically collected from real-world scenarios and contain identity-sensitive information such as faces, Fioresi et al. (Fioresi et al., 2023) proposed a privacy-preserving VAD framework named TeD-SPAD. They first anonymized video frames using a UNet to eliminate privacy information before using the I3D network to extract spatial-temporal features. Results showed that TeD-SPAD successfully prevented 32% of visual information leakage.
@@ 4.1.2. Multimodal Methods
Moreover, with the rise of Visual-Language Learning (VLL), particularly through the emergence of Large Language Models (LLMs), researchers have begun to leverage textual information to further enhance VAD models’ performance. Pre-trained Vision-Language Large Models (VLLMs) can describe appearance and motion information in videos without prior samples, embedding such text as prompts into visual representations. This integration represents a significant advancement in the fusion of video and text, which improves the model’s ability to express complex anomalies. The incorporation of textual information not only enhances generalizability and interpretability but also gives rise to new research tasks of practical value, such as video anomaly retrieval (Wu et al., 2024b) and open vocabulary VAD (Wu et al., 2024c), as illustrated in Sections 2.4 and 2.5.
In surveillance videos, it is often difficult to capture synchronized audio or text data, making unimodal methods the mainstream approach. However, with the rise of live streaming platforms and the film industry, multimodal methods will play a crucial role in online video moderation and content detection for TV shows, films, and animations. On the one hand, the context of such content is often highly complex and varied, with anomalies taking many forms, such as visual violence or non-compliant audio or text, making it difficult for unimodal methods relying solely on RGB sequences to handle effectively. On the other hand, these types of videos typically capture synchronized audio and provide subtitles or other multimodal data, allowing multimodal methods to function without additional data preparation costs.

### v3

@@ 3. Unsupervised Video Anomaly Detection
UVAD follows the general setup of semantic anomaly detection tasks, where only easily collectible regular instances are used to train models describing the normality of videos, detecting anomalies by measuring deviations between test samples and the learned model (Cheng et al., 2024a, b; Wang et al., 2021b). From traditional machine learning to deep representation learning (Yang et al., 2024a; Wang et al., 2024c; Yang et al., 2023b), UVAD has undergone multiple advancements in feature extraction and normality learning, leading VAD to become a key issue in the AD and CV communities. With its close ties to the AD community and its long-standing development history, UVAD has long been regarded by researchers as the mainstream research route of NSVAD algorithms. As a result, existing surveys (Nayak et al., 2021; Ramachandra et al., 2020) typically focus on reviewing UVAD literature, lacking in-depth discussion of emerging weakly supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Zaheer et al., 2022; Pang et al., 2020) routes, and overlooking novel tasks such as open-set and open-word detection. This article not only systematically reviews the latest developments in these new routes and tasks but also provides foundational knowledge and classic methods of UVAD in this section.
@@
Based on the means of normality learning and deviation calculation principles, existing deep learning methods are generally divided into three categories: distance-based (Saligrama and Chen, 2012; Dalal et al., 2006; Dalal and Triggs, 2005), probability-based (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), and reconstruction-based (Liu et al., 2024d, 2023g; Bao et al., 2022; Luo et al., 2017). From our perspective, distance-based UVAD methods are a more general form of probability-based and reconstruction-based methods because probability deviation and reconstruction error calculation are fundamentally just different distance measurement methods. Specifically, distance-based methods include using single-classifier learning of video sequence spatial-temporal representations in the deep feature space, such as OC-SVM and OC-NN (Ionescu et al., 2019; Sun et al., 2019). The drawback of such methods is that the trained models cannot be incrementally compatible with new data, leading to the need to retrain classifiers from scratch when new data is generated, such as in scenarios involving scene transitions. Another type of distance-based method is to use Gaussian mixture models (Wu et al., 2010) to model video feature normal vectors and measure deviations using Mahalanobis distance. In contrast, probability-based methods attempt to map the spatial-temporal representations of regular events into a probability framework and discriminate anomaly instances by measuring differences in probability distributions. Such methods tend to use traditional models such as Markov random fields (Benezeth et al., 2009) to build probability space. Deep learning-based attempts have encountered significant increases in computational costs and are noticeably slower in inference speeds.
@@
In fact, the most prevailing deep UVAD approach is reconstruction-based (Ramachandra et al., 2020). On the one hand, reconstruction-based methods (Liu et al., 2018b; Park et al., 2020; Liu et al., 2024d) aim to train models to represent the general spatial-temporal patterns of regular events through self-supervised proxy tasks, benefiting from advances in video self-supervised learning (Jing and Tian, 2020) and deep neural networks. On the other hand, such methods avoid complex mathematical computations, which are easy to implement and exhibit excellent performance, thus being widely praised by existing researchers. The premise of reconstruction-based methods is that generative models trained on massive normal samples can effectively infer the spatial-temporal patterns of regular events. For anomalous events, the performance of proxy tasks will significantly decrease, and the resulting error can be used as a quantitative basis for measuring deviations to calculate anomaly scores. Common generative models include deep autoencoders (Hasan et al., 2016; Chong and Tay, 2017; Liu et al., 2022a, e; Park et al., 2020), variational autoencoders (Fan et al., 2020; Yan et al., 2018), and generative adversarial networks (Chen et al., 2021; Nguyen et al., 2020; Liu et al., 2025c). Proxy tasks include reconstructing input sequences and predicting future frames, which belong to pixel-level image generation. However, most reconstruction-based methods only calculate frame-level errors as anomaly scores without performing spatial localization, as the spatial contribution of anomalies is typically not significant. Spatial localization can only serve as a quantitative visualization result.
@@ 3.1. Taxonomy and Advances
UVADGNLLPM Single-task: (Medel and Savakis, 2016), (Hasan et al., 2016), (Hu et al., 2016), (Smeureanu et al., 2017),(Luo et al., 2017),(Liu et al., 2018b),(Luo et al., 2021),(Singh and Pankajakshan, 2018),(Li et al., 2019), (Zhou et al., 2019b),(Nawaratne et al., 2019),(Ye et al., 2019),(Gong et al., 2019),(Park et al., 2020),(Luo et al., 2019),(Deepak et al., 2021),(Dong et al., 2020),(Yu et al., 2021),(Zhao et al., 2022),(Chen et al., 2020),(Wang et al., 2021a),(Nayak et al., 2023) Multi-tasks: (Chang et al., 2020),(Chang et al., 2021),(Li et al., 2021a),(Cai et al., 2021),(Liu et al., 2022a),(Liu et al., 2022e),(Zhao et al., 2017),(Xu et al., 2017),(Nguyen and Meunier, 2019),(Ravanbakhsh et al., 2019),(Zaheer et al., 2020) ST Patch: (Mahadevan et al., 2010),(Zhou et al., 2016),(Christiansen et al., 2016),(Sabokrou et al., 2017),(Wu et al., 2019),(Li et al., 2020),(Liu et al., 2021a),(Wang et al., 2018),(Roshtkhari and Levine, 2013) Object-level: (Georgescu et al., 2021b),(Georgescu et al., 2021a),(Hinami et al., 2017),(Ko and Sim, 2018),(Ionescu et al., 2019),(Yu et al., 2020),(Doshi and Yilmaz, 2021),(Liu et al., 2021d),(Chen et al., 2022),(Bao et al., 2022),(Liu et al., 2023a)
@@
GNL utilizes the entire video sequence as input, often requiring no additional preprocessing such as spatial-temporal cube partitioning or foreground object extraction, and employs end-to-end deep neural networks to directly learn video ecologies (Huang et al., 2022; Gong et al., 2019; Park et al., 2020; Chang et al., 2020; Liu et al., 2024d, 2023f). Over the years, researchers have believed that videos possess two informational dimensions, namely spatial and temporal, corresponding to appearance and motion, requiring different approaches to capture their normality. Therefore, within our UVAD taxonomy, GNL is further subdivided into Single-Proxy Task (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020) and Multi-Proxy Task methods (Chang et al., 2020; Liu et al., 2022e, 2024d). In contrast, LPM methods argue that video data contain a plethora of redundant information tantamount to clues related to normality. Thus, they opt to use spatial-temporal cubes (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a) or foreground objects (Bao et al., 2022; Liu et al., 2023a) containing dense and effective information as network inputs instead of the entire video sequence, focusing on prototype feature learning of local patch. We classify such methods into spatial-temporal Patch-based and Object-driven methods.
@@ 3.1.1. Global Normality Learning
The earliest methods did not distinguish between spatial and temporal information, typically using only RGB videos as input and employing a single self-supervised proxy task (e.g., sequence reconstruction or future frame prediction (Zhao et al., 2024)) to broadly learn spatial-temporal normality. Generally, methods based on a single proxy task focus on designing more efficient single-stream end-to-end deep structures. Recent efforts include the introduction of 3D convolutional networks (Gong et al., 2019) and convolutional long short-term memory networks (Luo et al., 2017) to enhance the representation capability of spatial-temporal features. Subsequent researchers discovered that spatial and temporal information have different characteristics, with the former focusing on local pixel inference and the latter on modeling global dynamics. Moreover, the addition of proxy tasks such as reconstruction and prediction losses often brings additional performance gains without significantly increasing training costs. Thus, they proposed introducing additional proxy tasks within the GNL framework. In addition to the separation of spatial-temporal normality learning (Liu et al., 2022e; Chang et al., 2021) and simple proxy task stacking (Zhao et al., 2017; Chen et al., 2020), recent work has also explored novel tasks such as appearance-motion consistency (Cai et al., 2021), spatial-temporal coherence (Cheng et al., 2023a; Ning et al., 2024), and correlation (Zhao et al., 2022).
@@ 4. Weakly-supervised Video Anomaly Detection
Inspired by multiple instance learning (MIL) (Herrera et al., 2016), WsVAD organizes videos into bags containing several instances. All segment instances from the same bag share a video-level label (Huang et al., 2024a). Under this setting, $Y=0$ indicates negative bags while $Y=1$ indicates positive ones containing at least one anomalous instance. WsVAD strikes a balance between performance and data preparation cost, bridging the gap between UVAD with unreliable results and supervised learning with fine labels. Specifically, unlike unsupervised methods that train models only on regular events and simply consider all unseen samples as anomalies, WsVAD is trained with positive samples, enabling it to effectively understand the inherent differences between normal and anomalous instances. Consequently, existing research indicates that WsVAD yields more reliable results and outperforms UVAD schemes with lower false alarm rates in real-world tasks, such as crime behavior identification and traffic accident detection (Liu et al., 2023d). Annotating long video sequences frame by frame for supervised learning is often impractical. For instance, the training set of the UCF-Crime dataset includes 1,610 long videos containing 13,741,393 frames in total. In contrast, WsVAD only requires coarse semantic video-level annotations, with each video needing only a discrete binary label. Various types of anomalies are roughly labeled as 1, requiring minimal expert involvement in data annotation and significantly reducing labor costs compared to fine-grained labeling. Thus, although WsVAD no longer adheres to the open-world task setting of the AD community and can only detect specific anomalies, it has garnered widespread attention in recent years due to its stable performance and reliable results. Subsequent works typically adopt the MIL regression task setting proposed by Sultani et al. (Sultani et al., 2018) and utilize their concurrently open-sourced UCF-Crime dataset as a benchmark. To further validate the generalization ability of weakly-supervised models, some researchers relocate positive samples from the test set of UVAD datasets to the training sets and provide video-level labels, proposing reconfigured datasets such as UCSD Ped2 and ShanghaiTech Weakly for WsVAD validation.
@@ 4.1. Taxonomy and Advances
The superior performance of WsVAD on real-world videos has inspired researchers to develop VAD models tailored to the complexities of real-life scenarios. In 2020, Wu et al. (Wu et al., 2020) extended the research scope of VAD from single-modal video pattern analysis to multimodal learning to leverage heterogeneous data in real-world scenes for enhanced anomaly event detection. Their collected XD-violence dataset is the first multimodal VAD benchmark, comprising RGB images and audio modalities, focusing on violence behavior detection in complex scenes. This work not only provides the first multimodal VAD dataset and solution but also motivates researchers in the community to explore anomaly clues from multimodal data such as audio and text, leading to a new wave of VAD development: multimodal WsVAD. In this section, we comprehensively review the latest advancements in both single-modal and multimodal WsVAD, and illustrate the foundational knowledge and specific implementations required for WsVAD using the Multiple Instance Regression (MIR) framework and the Local-Global Network (HL-Net) as examples.
@@ 4.1.1. Unimodal Methods
Unimodal WsVAD (UWsVAD) methods focus on extracting anomaly-related cues from RGB image sequences (Liu et al., 2022f). These methods generally follow a three-step process: 1) preprocessing videos into several non-overlapping segments, 2) extracting spatial-temporal features using models like Convolutional 3D Networks (Tran et al., 2015) and Inflated 3D Networks (Carreira and Zisserman, 2017), and 3) computing anomaly scores using a multi-instance ranking loss to differentiate between normal and anomalous instances. The Multi-Instance Ranking (MIR) framework (Sultani et al., 2018) first introduced MIL to WsVAD with a focus on predicting higher anomaly scores for anomalous segments while minimizing score fluctuations. For a more detailed description of the existing UWsVAD methods, please refer to Section 1.2.1 of Supplement. Moreover, given that VsVAD datasets are typically collected from real-world scenarios and contain identity-sensitive information such as faces, Fioresi et al. (Fioresi et al., 2023) proposed a privacy-preserving VAD framework named TeD-SPAD. They first anonymized video frames using a UNet to eliminate privacy information before using the I3D network to extract spatial-temporal features. Results showed that TeD-SPAD successfully prevented 32% of visual information leakage.
@@ 4.1.2. Multimodal Methods
Moreover, with the rise of Visual-Language Learning (VLL), particularly through the emergence of LLMs, researchers have begun to leverage textual information to further enhance VAD models’ performance. Pre-trained Vision Vision LLMs can describe appearance and motion information in videos without prior samples, embedding such text as prompts into visual representations. This integration represents a significant advancement in the fusion of video and text, which improves the model’s ability to express complex anomalies. The incorporation of textual information not only enhances generalizability and interpretability but also gives rise to new research tasks of practical value, such as video anomaly retrieval (Wu et al., 2024b) and open vocabulary VAD (Wu et al., 2024c), as illustrated in Sections 2.4.
Specifically, VLL models (Yang et al., 2024c) can provide accurate textual descriptions of video frames, enhancing the semantic mining capability of existing vision-based VAD models. For instance, Chen et al. (Chen et al., 2023) used a language model-based captioning network to obtain textual descriptions of video sequences, which, after being embedded in a text embedding network, were fused with visual features as inputs to the anomaly detector. Their proposed Text Empowered Video Anomaly Detection (TEVAD) efficiently captures abstract semantics of anomaly events and enhances the interpretability of VAD models. Pu et al. (Pu et al., 2024) introduced a Prompt-Enhanced Learning (PEL) module, using knowledge-based prompts to incorporate semantic priors, improving the discriminative power of visual features in weakly supervised VAD, while ensuring separability between anomalous subcategories. Wu et al. (Wu et al., 2024b) proposed VAR, which efficiently detects specific anomalies based on cross-modal learning (e.g., language descriptions and synchronized audio). They designed an Anomaly-Led Alignment Network (ALAN) using BERT (Kenton and Toutanova, 2019) to process text information and incorporate a pretext task to enhance semantic alignment between video-text fine-grained representations.
In addition, visual-language associations can serve as effective cues for detecting video anomalies, and the pre-learned visual-text consistency in large VLL models can be efficiently transferred to VAD. Kim et al. (Kim et al., 2023) used large language models to generate textual descriptions of video frames and detected anomalous frames by calculating the cosine similarity between input frames and their textual descriptions using CLIP. Text Prompt with Normality Guidance (Yang et al., 2024b) leverages the language-visual knowledge of the CLIP model to align video frames with textual descriptions of events, generating more accurate pseudo-labels for WsVAD, thus improving model performance.
In surveillance videos, it is often difficult to capture synchronized audio or text data, making unimodal methods the mainstream approach. However, with the rise of live streaming platforms and the film industry, multimodal methods will play a crucial role in online video moderation and content detection for TV shows, films, and animations. On the one hand, the context of such content is often highly varied, with anomalies taking many forms, such as visual violence or non-compliant audio or text, making it difficult for unimodal methods relying solely on RGB sequences to handle effectively. On the other hand, these types of videos typically capture synchronized audio and provide subtitles or other multimodal data, allowing multimodal methods to function without additional data preparation costs.